20. Ram is now 15 years older than his younger brother Govind. Ten years from now, Ram will be twice as old as Govind. Find the present ages of each.

Answer by dkppathak(439)

let govind is x yrs
ram is x+15 yrs
10 yrs from now
govind =x+10
ram is x+25
as per given conditions
x+25=2(x+10)
x+25=2x+20
x=5 yrs
govind is 5yrs and Ram is 5 yrs